Scary Store Ltd sells Halloween costumes every year. Each costume costs the store $35, and retails for $110. The costumes can be sold for $10 once Halloween is over. Based on past demands, the store managed predicts his demand to be 240 costumes with a standard deviation of 80 costumes. How many costumes should the store manager order for this year’s Halloween?


1
Expert's answer
2020-08-17T19:07:06-0400

If I sell a purchased suit, I will earn 110-35 = 75

If I don't sell the suit, I will lose 35-10 = 25

We get a ratio of 3: 1

Based on the possible variation, the minimum number of purchased suits is 240-80 = 160 and the maximum number of possible purchased suits is 240 + 80 = 320

If we take 320 for 100% and 160 for 0% and divide it in a ratio of 3: 1, then we get that it is necessary to buy 160+ (320-160) * 3/4 = 280